# Continuum Attention for Neural Operators

### v1

@@ 6.2.3 Kolmogorov Flow
where $u$ denotes the velocity, $p$ the pressure and $\nu$ the kinematic viscosity. The particular choice of forcing function $\sin(ny)$ leads to a particular example of a Kolmogorov flow. We equip the domain $[0,2\pi]$ with periodic boundary conditions. We assume
@@
The dot on $L^{2}$ denotes the space of spatially mean-zero functions. The vorticity is defined as $w=(\nabla\times u)\widehat{z}$ and the stream function $f$ as the solution to the Poisson equation $-\Delta f=w$. Existence of the semigroup $S_{t}:\mathcal{U}\to\mathcal{U}$ is shown in Temam (2012, Theorem 2.1). We generate the data by solving (KF) in vorticity-streamfunction form by applying the pseudo-spectral split step method from Chandler and Kerswell (2013). In our experimental set-up $n=4$ and $\nu=1/70$ are chosen. Random initial conditions are sampled from the Gaussian measure $\mathcal{N}(0,C)$ where the covariance operator is given by
@@
where the Laplacian is equipped with periodic boundary conditions on $[0,2\pi]^{2}$, and viewed as acting between spaces of spatially mean-zero functions. The input output-data pairs for the Kolmogorov flow experiment are given by
@@
In our experimental setting, we set $T=11$ and $\Delta t=0.1$. We train the FNO, ViT neural operator and Fourier attention neural operator architectures on this problem using $9000$ independent samples of resolution $416\times 416$, $500$ samples for validation and $500$ samples for testing. The Fourier neural operator is parametrized by $12$ Fourier modes in each dimension and channel width of $128$; the model is trained using a batch size of $4$ and learning rate of $10^{-3}$. Both the ViT neural operator and Fourier attention neural operator are parametrized by $128$ channels in the transformer encoder. The integral operator in the ViT NO is parametrized by $12$ Fourier modes in each dimension while the integral operators appearing in the Fourier attention neural operator are parametrized by $9$ Fourier modes in each dimension. Both architectures employ a final spectral convolution layer (Remark 21) that is parametrized by $64$ Fourier modes in each dimension. Both of these neural operators are trained using a batch size of $2$ and learning rate of $10^{-3}$. On this problem we train all the neural operators with the relative $H^{1}$ loss. Figure 15 displays the test results using the Fourier attention neural operators.
@@
A few additional considerations are in order. The mode truncation employed in the Fourier neural operator make the architecture computationally efficient but yields an over-smoothing effect that is not suitable for problems at high resolutions, where high frequency detail is prominent. Attention-based neural operators offer a possible solution to this issue, as demonstrated by the performance of the transformer neural operators proposed when applied to the PDE operator learning problems considered, which involve rough diffusion coefficients and rough initial conditions. On the other hand, patching leads to more efficient attention-based neural operators, but also introduces possible discontinuities at patch-intersections. This issue has been observed to be resolved in the large data regime. The residual spectral convolution layers employed ameliorate the discontinuities that are visible in pointwise error plots, but have a negligible effect on relative $L^{2}$ error in training and testing. This property constitutes an inherent limitation of patching and highlights the need for more suitable efficient attention mechanisms. Furthermore, it is observed that in some experimental settings the patching-based neural operators exhibit worse stability to changing resolution than FNO. This is likely due to the effect of patching which may then be further amplified for FANO due to the application of a higher number of FFT applications to non-periodic domains. Such issues could potentially be overcome using periodic extensions for Fourier-based parametrizations or other kinds of parametrizations for the nonlocal operations, or could be further improved by fine-tuning.
@@ 7 Conclusions
In this work we have introduced a continuum formulation of the attention mechanism from the seminal work Vaswani et al. (2017). Our continuum formulation can be used to design transformer neural operators. Indeed, this continuum formulation leads to discretization invariant implementations of attention and schemes that exhibit zero-shot generalization to different resolutions. In the first result of its kind for transformers, the resulting neural operator architecture mapping between infinite-dimensional spaces of functions is shown to be a universal approximator of continuous functions and functions of Sobolev regularity defined over a compact domain. We extend the continuum formulation to patched attention, which makes it possible to design efficient transformer neural operators. To this end, we introduce a neural operator analogue of the vision transformer Dosovitskiy et al. (2021) and a more expressive architecture, the Fourier attention neural operator. Through a cost-accuracy analysis, we demonstrate the power of the methodology in the context of high-resolution operator learning problems.
@@ B.2 Vision Transformer Neural Operator
for any $u\in\mathbbm{R}^{P\times p_{1}\times\ldots\times p_{d}\times(d_{u}+d)}$, where we have made explicit the dependence on $\theta\in\Theta$ to highlight the learnable components in the implementation. The FFT is computed on the $p_{1}\times\ldots\times p_{d}$ dimensions so that the resulting tensor is of the form $\mathcal{F}{u}\in\mathbbm{C}^{P\times p_{1}\times\ldots\times p_{d}\times(d_{u}+d)}$. Since ${u}$ is convolved with a function that possess only $\bigl{|}Z_{k_{\textrm{max}}}\bigr{|}$ Fourier modes, the higher Fourier modes may be truncated so that one may consider $\mathcal{F}{u}\in\mathbbm{C}^{P\times 2k_{\textrm{max},1}\times\ldots\times 2k_{\textrm{max},d}\times(d_{u}+d)}$. The learnable tensor is $R_{\mathsf{W}_{\textrm{in}}}\in\mathbbm{C}^{2k_{\textrm{max},1}\times\ldots\times 2k_{\textrm{max},d}\times d_{\textrm{model}}\times(d_{u}+d)}$ and the operation in Fourier space is defined as
@@
for $p\in\llbracket 1,P\rrbracket,\,1\leq i_{j}\leq 2k_{\textrm{max},j},\,1\leq\ell\leq d_{\textrm{model}}$. The inverse FFT is then applied along the dimensions $2k_{\textrm{max},1},\ldots,2k_{\textrm{max},d}$. The resulting tensor defined as ${v}^{(0)}\in\mathbbm{R}^{{P\times p_{1}\times\ldots\times p_{d}\times d_{\textrm{model}}}}$ is then fed through the transformer encoder $\mathsf{E}_{L}:\mathbbm{R}^{{P\times p_{1}\times\ldots\times p_{d}\times d_{\textrm{model}}}}\to\mathbbm{R}^{{P\times p_{1}\times\ldots\times p_{d}\times d_{\textrm{model}}}}$.
@@
for any $v\in\mathbbm{R}^{n_{1}\times\ldots\times n_{d}\times d_{\textrm{model}}}$. Following Remark 21, an additional convolution in frequency space may be applied in residual form, where application of the nonlocal integral operator follows Definition 20 and the subsequent discussion. To this end, the operation in frequency space is defined similarly to (61) and the learnable tensor is of the form $R_{\mathsf{W}_{\textrm{S}}}\in\mathbbm{C}^{k_{\textrm{max},1}\times\ldots\times k_{\textrm{max},d}\times d_{z}\times d_{z}}.$
@@ B.3 Fourier Attention Neural Operator
for $h=1,\ldots,H$ and any $v\in\mathbbm{R}^{p_{1}\times\ldots\times p_{d}\times d_{\textrm{model}}}$, where we have made explicit the dependence on $\theta\in\Theta$ to highlight the learnable components in the implementation. In (68) the FFT is computed on the $p_{1}\times\ldots\times p_{d}$ so that the resulting tensor $\mathcal{F}{v}\in\mathbbm{C}^{p_{1}\times\ldots\times p_{d}\times d_{\textrm{model}}}$. Since ${v}$ is a function that possess only $\bigl{|}Z_{k_{\textrm{max}}}\bigr{|}$ Fourier modes, the higher Fourier modes may be truncated so that one may consider $\mathcal{F}{v}\in\mathbbm{C}^{2k_{\textrm{max},1}\times\ldots\times 2k_{\textrm{max},d}\times d_{\textrm{model}}}$. The learnable tensors are of dimension $R_{\mathsf{Q}_{h}},R_{\mathsf{K}_{h}},R_{\mathsf{V}_{h}}\in\mathbbm{C}^{2k_{\textrm{max},1}\times\ldots\times 2k_{\textrm{max},d}\times d_{K}\times d_{\textrm{model}}}$ for $h=1,\ldots,H$, and the operation in Fourier space is defined as
@@
for any $v\in\mathbbm{R}^{p_{1}\times\ldots\times p_{d}\times d_{\textrm{model}}}$ and  $\,1\leq i_{j}\leq 2k_{\textrm{max},j},\,1\leq\ell\leq d_{K}$. The inverse FFT is computed along the dimensions $k_{\textrm{max},1},\ldots,k_{\textrm{max},d}$ so that $\mathsf{Q}_{h}{v},\mathsf{K}_{h}{v},\mathsf{V}_{h}{v}\in\mathbbm{R}^{p_{1}\times\ldots\times p_{d}\times d_{K}}$ for $h=1,\ldots,H$. We recall that in this case the self-attention operator $\mathsf{A}$ is defined according to Definition 14. While the expectation is computed with respect to a probability measure on the patch sequence, hence inducing a discrete probability mass function, the discretization of the domain introduces the need to approximate the $L^{2}$ inner product appearing in Definition 13. We resort to an approximation of the integral given by (62) and (63). The multi-head attention operator is thus defined by the application of the linear transformation $W_{\textrm{MultiHead}}^{\top}\in\mathbbm{R}^{Hd_{K}\times d_{\textrm{model}}}$ to the concatenation of $H\in\mathbbm{N}$ self-attention operations. The concatenation of the outputs of the $H\in\mathbbm{N}$ self-attention operations results in a $\mathbbm{R}^{P\times p_{1}\times\ldots\times p_{d}\times Hd_{K}}$ tensor, so that
@@
A reshaping operator $\mathsf{S}^{\prime}_{\textrm{reshape}}$ to the tensor resulting from the transformer encoder iteration as in Equation 66. A pointwise linear transformation $W_{\textrm{out}}^{\top}\in\mathbbm{R}^{d_{\textrm{model}}\times d_{z}}$ is then applied as in Equation 67. Again, following Remark 21, an additional convolution in frequency space may be applied in residual form, where application of the nonlocal integral operator follows Definition 20 and the subsequent discussion. To this end, the operation in frequency space is defined similarly to (61) and the learnable tensor is of the form $R_{\mathsf{W}_{\textrm{S}}}\in\mathbbm{C}^{k_{\textrm{max},1}\times\ldots\times k_{\textrm{max},d}\times d_{z}\times d_{z}}.$
@@ C.1 Parameter Complexity
Table 4: Parameter complexity for the three proposed transformer neural operators, as implemented in practice, compared to the Fourier neural operator (Li et al., 2021). We note that for our implementations, $d_{\textrm{FNO}}$, which is the latent dimension of the two-layer lifting and projection MLP in FNO, is fixed to be $128$. Parameters arising from possible bias terms, parameters arising from skip connections in FNO, as well as parameters arising from a final convolution employed for the patching architectures are omitted for ease of presentation.
@@ C.2 Evaluation Cost
FNO $&2N(d_{u}+d+2d_{\textrm{model}})d_{\textrm{FNO}}+2Nd_{\textrm{FNO}}d_{z}\\ &+4\bigl{(}15d_{\textrm{model}}N\log(\sqrt{N})+k_{\textrm{max}}\cdot(2d^{2}_{\textrm{model}}-d_{\textrm{model}})\\ &+2Nd^{2}_{\textrm{model}}\bigr{)}\end{aligned}$
Transformer NO $&2N(d_{u}+d)d_{\textrm{model}}+2Nd_{\textrm{model}}d_{z}+6\bigl{(}8d^{2}_{\textrm{model}}N\\ &+2N^{2}d_{\textrm{model}}+4d^{2}_{\textrm{model}}N\bigl{)}\end{aligned}$
ViT NO $&(d_{u}+d+d_{\textrm{model}})\cdot 15N\log(\sqrt{N/P})/2\\ &+Pk_{\textrm{max}}d_{\textrm{model}}\bigl{(}2(d_{u}+d)-1\bigr{)}+2Nd_{\textrm{model}}d_{z}\\ &+6\Bigl{(}8d^{2}_{\textrm{model}}N+2d_{\textrm{model}}NP+4d^{2}_{\textrm{model}}N\Bigr{)}\\ &+15d_{z}N\log(\sqrt{N})+k^{\prime}_{\textrm{max}}(2d^{2}_{z}-d_{z})\end{aligned}$
FANO $&2N(d_{u}+d)d_{\textrm{model}}+2Nd_{\textrm{model}}d_{z}\\ &+6\Bigl{(}45d_{\textrm{model}}N\log(\sqrt{N/P})+3k_{\textrm{max}}P\bigl{(}2d^{2}_{\textrm{model}}-d_{\textrm{model}}\bigr{)}\\ &+2Nd^{2}_{\textrm{model}}+2d_{\textrm{model}}NP+4Nd^{2}_{\textrm{model}}\Bigr{)}\\ &+15d_{z}N\log(\sqrt{N})+k^{\prime}_{\textrm{max}}(2d^{2}_{z}-d_{z})\end{aligned}$
Table 5: Evaluation cost (in FLOPS) for the proposed transformer neural operators, as implemented in practice, compared to FNO (Li et al., 2021). As done in Kaplan et al. (2020) and for ease of presentation, the additional evaluation cost arising from nonlinear activations, the softmax operator and normalizations are omitted. Furthermore, we omit the cost arising from skip connections. Indeed these do not affect the scaling. We further note that for our implementations, $d_{\textrm{FNO}}$, which is the latent dimension of the two-layer lifting and projection MLP in FNO, is fixed to be $128$.

### v2

@@ 6.2.3 Kolmogorov Flow
where $u$ denotes the velocity, $p$ the pressure and $\nu$ the kinematic viscosity. The particular choice of forcing function $\sin(ny)$ leads to a particular example of a Kolmogorov flow. We equip the domain $0,2\pi]$ with periodic boundary conditions. We assume
@@
The dot on $L^{2}$ denotes the space of spatially mean-zero functions. The vorticity is defined as $w=(\nabla\times u)\widehat{z}$ and the stream function $f$ as the solution to the Poisson equation $\Delta f=w$. Existence of the semigroup $S_{t}:\mathcal{U}\to\mathcal{U}$ is shown in Temam (2012, Theorem 2.1). We generate the data by solving (KF) in vorticity-streamfunction form by applying the pseudo-spectral split step method from Chandler and Kerswell (2013). In our experimental set-up $n=4$ and $\nu=1/70$ are chosen. Random initial conditions are sampled from the Gaussian measure $\mathcal{N}(0,C)$ where the covariance operator is given by
@@
where the Laplacian is equipped with periodic boundary conditions on $0,2\pi]^{2}$, and viewed as acting between spaces of spatially mean-zero functions. The input output-data pairs for the Kolmogorov flow experiment are given by
@@
In our experimental setting, we set $T=11$ and $\Delta t=0.1$. We train the AFNO, FNO, ViT neural operator and Fourier attention neural operator architectures on this problem using $9000$ independent samples of resolution $416\times 416$, $500$ samples for validation and $500$ samples for testing. The Fourier neural operator is parametrized by $12$ Fourier modes in each dimension and channel width of $128$; the model is trained using a batch size of $4$ and learning rate of $10^{-3}$. The AFNO is instantiated with $4$ layers of $128$ hidden channels, and does not employ patching; it is trained using a learning rate of $10^{-3}$ and batch size of $1$. Both the ViT neural operator and Fourier attention neural operator are parametrized by $128$ channels in the transformer encoder. The integral operator in the ViT NO is parametrized by $12$ Fourier modes in each dimension while the integral operators appearing in the Fourier attention neural operator are parametrized by $9$ Fourier modes in each dimension. Both architectures employ a final spectral convolution layer (Remark 21) that is parametrized by $64$ Fourier modes in each dimension. Both of these neural operators are trained using a batch size of 22 and learning rate of $10^{-3}$. On this problem we train all the neural operators with the relative $H^{1}$ loss. Figure 16 displays the test results using the Fourier attention neural operators.
@@
A few additional considerations are in order. The mode truncation employed in the Fourier neural operator make the architecture computationally efficient but yields an over-smoothing effect that is not suitable for problems at high resolutions, where high frequency detail is prominent. Attention-based neural operators offer a possible solution to this issue, as demonstrated by the performance of the transformer neural operators proposed when applied to the PDE operator learning problems considered, which involve rough diffusion coefficients and rough initial conditions. On the other hand, patching leads to more efficient attention-based neural operators, but also introduces possible discontinuities at patch-intersections. This issue has been observed to be resolved in the large data regime. Furthermore, it is possible to employ problem specific smoothing operators as the one introduced in Remark 21. In Figure 18 we display the result of applying a smoothing operator as the last layer in the training of the FANO architecture in the context of the Kolmogorov flow problem. The smoothing via the inverse of the negative Laplacian is applied using the Fourier basis, given the periodic boundary conditions. Here, we make the specific choice $\epsilon=10^{-3}$ and $\alpha=1.001$, in the context of Remark 21. Such an approach is effective at reducing the discontinuities introduced by patching and reducing the evaluation error, but is problem dependent and requires knowledge of boundary conditions. Such discontinuity issues constitute an inherent limitation of patching and highlights the need for more suitable efficient attention mechanisms.
Figure 18: The panel displays the result of the application of the Fourier attention neural operator employing a final smoothing layer on the Kolmogorov flow experiment for the median and maximum relative $L^{2}$ error samples. The first row displays the sample from the test set of the same resolution as the training set yielding the median relative $L^{2}$ error. The second row on the other hand displays the sample yielding the maximum relative $L^{2}$ error. The first column of the panel displays the input vorticity $w(x,T)$ of the relevant sample. The second column shows the absolute difference between the true vorticity $w(x,T+\Delta t)$ and the input vorticity $w(x,T)$, while the third column displays the absolute difference between the predicted vorticity at time $T+\Delta t$ and the input $w(x,T)$. The last column shows the pointwise absolute error (in log-scale) between the prediction and truth.
Furthermore, it is observed that in some experimental settings the patching-based neural operators exhibit worse stability to changing resolution than FNO. This is likely due to the effect of patching which may then be further amplified for FANO due to the application of a higher number of FFT(s) to non-periodic domains. In the context of the FANO architecture, such issues can be ameliorated using periodic extensions for Fourier-based parametrizations. In Figure 19 we demonstrate the discretization invariance of the standard FANO compared to a FANO architecture that involves building a periodic extension of each patch. The Fourier integral operator is applied to each extended patch and the extension is then removed from the output. Such a technique leads to an increase in complexity and runtime but yields more stable discretization invariance. Other remedies could potentially include other kinds of parametrizations for the nonlocal operations, or fine-tuning at different resolutions.
Figure 19: The panel displays the performance in relative $L^{2}$ errors (in log-scale) of the Fourier attention neural operator and Fourier attention neural operator with periodic extensions of patches when applied to Kolomgorov flow test sets of different resolutions at inference time (without retraining). All architectures are trained on data that is of resolution $416\times 416$.
@@ 7 Conclusions
In this work we have introduced a continuum formulation of the attention mechanism from the seminal work Vaswani et al. (2017). Our continuum formulation can be used to design transformer neural operators. Indeed, this continuum formulation leads to discretization invariant implementations of attention and schemes that exhibit zero-shot generalization to different resolutions. In the first result of its kind for transformers, with a slight modification to the architecture implemented in practice, the resulting neural operator architecture mapping between infinite-dimensional spaces of functions is shown to be a universal approximator of continuous functions and functions of Sobolev regularity defined over a compact domain. We extend the continuum formulation to patched attention, which makes it possible to design efficient transformer neural operators. To this end, we introduce a neural operator analogue of the vision transformer Dosovitskiy et al. (2021) and a more expressive architecture, the Fourier attention neural operator. Through a cost-accuracy analysis, we demonstrate the power of the methodology in the context of a range of operator learning problems. In the following we highlight potential avenues for further work.
  1. 1.
    It is of interest to extend the analysis of Theorems 6 and 12 to quantify the norm-dependent rates of convergence of discretized attention to its continuum limit. Furthermore, obtaining the results over i.i.d. samples from arbitrary strictly positive probability distributions would be of interest.
  2. 2.
    The Fourier attention neural operator architecture uses operator parametrizations for $\mathsf{Q},\mathsf{K},\mathsf{V}$. Other design choices for these operators not employing the Fourier basis are possible, for example using the wavelet basis as in Tripura and Chakraborty (2023) or using a multigrid approach as in He et al. (2024). An investigation of other attention neural operator architectures arising from this consideration would be of interest.
  3. 3.
    It is of interest to investigate the discontinuity issues arising from patching. In particular, it is desirable to design schemes that impose continuity amongst these throughout the architecture.
  4. 4.
    Deriving universal approximation theorems for the patch-based transformer neural operators would be of great theoretical interest.
@@ B.2 Vision Transformer Neural Operator
for any $u\in\mathbbm{R}^{P\times p_{1}\times\ldots\times p_{d}\times(d_{u}+d)}$, where we have made explicit the dependence on $\theta\in\Theta$ to highlight the learnable components in the implementation. The FFT is computed on the $p_{1}\times\ldots\times p_{d}$ dimensions so that the resulting tensor is of the form $\mathcal{F}{u}\in\mathbbm{C}^{P\times p_{1}\times\ldots\times p_{d}\times(d_{u}+d)}$. Since u{u} is convolved with a function that possess only $\bigl{|}Z_{k_{\textrm{max}}}\bigr{|}$ Fourier modes, the higher Fourier modes may be truncated so that one may consider $\mathcal{F}{u}\in\mathbbm{C}^{P\times(2k_{\textrm{max},1}+1)\times\ldots\times(2k_{\textrm{max},d}+1)\times(d_{u}+d)}$. The learnable tensor is $R_{\mathsf{W}_{\textrm{in}}}\in\mathbbm{C}^{(2k_{\textrm{max},1}+1)\times\ldots\times(2k_{\textrm{max},d}+1)\times d_{\textrm{model}}\times(d_{u}+d)}$ and the operation in Fourier space is defined as
@@
for $p\in\llbracket 1,P\rrbracket,\,1\leq i_{j}\leq 2k_{\textrm{max},j}+1,\,1\leq\ell\leq d_{\textrm{model}}$. The inverse FFT is then applied along the dimensions $2k_{\textrm{max},1}+1,\ldots,2k_{\textrm{max},d}+1$. The resulting tensor defined as ${v}^{(0)}\in\mathbbm{R}^{{P\times p_{1}\times\ldots\times p_{d}\times d_{\textrm{model}}}}$ is then fed through the transformer encoder $\mathsf{E}_{L}:\mathbbm{R}^{{P\times p_{1}\times\ldots\times p_{d}\times d_{\textrm{model}}}}\to\mathbbm{R}^{{P\times p_{1}\times\ldots\times p_{d}\times d_{\textrm{model}}}}$.
@@
for any $v\in\mathbbm{R}^{n_{1}\times\ldots\times n_{d}\times d_{\textrm{model}}}$.
@@ B.3 Fourier Attention Neural Operator
for $h=1,\ldots,H$ and any $v\in\mathbbm{R}^{p_{1}\times\ldots\times p_{d}\times d_{\textrm{model}}}$, where we have made explicit the dependence on $\theta\in\Theta$ to highlight the learnable components in the implementation. In (68) the FFT is computed on the $p_{1}\times\ldots\times p_{d}$ so that the resulting tensor $\mathcal{F}{v}\in\mathbbm{C}^{p_{1}\times\ldots\times p_{d}\times d_{\textrm{model}}}$. Since v{v} is a function that possess only $\bigl{|}Z_{k_{\textrm{max}}}\bigr{|}$ Fourier modes, the higher Fourier modes may be truncated so that one may consider $\mathcal{F}{v}\in\mathbbm{C}^{(2k_{\textrm{max},1}+1)\times\ldots\times(2k_{\textrm{max},d}+1)\times d_{\textrm{model}}}$. The learnable tensors are of dimension $R_{\mathsf{Q}_{h}},R_{\mathsf{K}_{h}},R_{\mathsf{V}_{h}}\in\mathbbm{C}^{(2k_{\textrm{max},1}+1)\times\ldots\times(2k_{\textrm{max},d}+1)\times d_{K}\times d_{\textrm{model}}}$ for $h=1,\ldots,H$, and the operation in Fourier space is defined as
@@
for any $v\in\mathbbm{R}^{p_{1}\times\ldots\times p_{d}\times d_{\textrm{model}}}$ and  $1\leq i_{j}\leq 2k_{\textrm{max},j}+1,\,1\leq\ell\leq d_{K}$. The inverse FFT is computed along the dimensions $2k_{\textrm{max},1}+1,\ldots,2k_{\textrm{max},d}+1$ so that $\mathsf{Q}_{h}{v},\mathsf{K}_{h}{v},\mathsf{V}_{h}{v}\in\mathbbm{R}^{p_{1}\times\ldots\times p_{d}\times d_{K}}$ for $h=1,\ldots,H$. We recall that in this case the self-attention operator $\mathsf{A}$ is defined according to Definition 14. While the expectation is computed with respect to a probability measure on the patch sequence, hence inducing a discrete probability mass function, the discretization of the domain introduces the need to approximate the $L^{2}$ inner product appearing in Definition 13. We resort to an approximation of the integral given by (62) and (63). The multi-head attention operator is thus defined by the application of the linear transformation $W_{\textrm{MultiHead}}^{\top}\in\mathbbm{R}^{Hd_{K}\times d_{\textrm{model}}}$ to the concatenation of $H\in\mathbbm{N}$ self-attention operations. The concatenation of the outputs of the $H\in\mathbbm{N}$ self-attention operations results in a $\mathbbm{R}^{P\times p_{1}\times\ldots\times p_{d}\times Hd_{K}}$ tensor, so that
@@
A reshaping operator $\mathsf{S}^{\prime}_{\textrm{reshape}}$ to the tensor resulting from the transformer encoder iteration as in Equation 66. A pointwise linear transformation $W_{\textrm{out}}^{\top}\in\mathbbm{R}^{d_{\textrm{model}}\times d_{z}}$ is then applied as in Equation 67.
@@ C.1 Parameter Complexity
AFNO $(d_{u}+d)\cdot d_{\textrm{model}}+N^{2}d_{\textrm{model}}+d_{\textrm{model}}d_{z}+4\cdot(8+4/b)\cdot d^{2}_{\textrm{model}}$
@@
Table 4: Parameter complexity for the three proposed transformer neural operators, as implemented in practice, compared to the Fourier neural operator (Li et al., 2021). We note that for our implementations, $d_{\textrm{FNO}}$, which is the latent dimension of the two-layer lifting and projection MLP in FNO, is fixed to be $128$, while the hyperparameter $b$ in the AFNO is fixed to be $16$. Parameters arising from possible bias terms, parameters arising from skip connections in FNO and parameters arising from normalizations in AFNO are omitted for ease of presentation.
@@ C.2 Evaluation Cost
FNO $\begin{aligned} &2N(d_{u}+d+2d_{\textrm{model}})d_{\textrm{FNO}}+2Nd_{\textrm{FNO}}d_{z}\\ &+4\bigl{(}15d_{\textrm{model}}N\log(\sqrt{N})+k_{\textrm{max}}\cdot(2d^{2}_{\textrm{model}}-d_{\textrm{model}})\\ &+2Nd^{2}_{\textrm{model}}\bigr{)}\end{aligned}$
Transformer NO $\begin{aligned} &2N(d_{u}+d)d_{\textrm{model}}+2Nd_{\textrm{model}}d_{z}+6\bigl{(}8d^{2}_{\textrm{model}}N\\ &+2N^{2}d_{\textrm{model}}+4d^{2}_{\textrm{model}}N\bigl{)}\end{aligned}$
[…]
Table 5: Evaluation cost (in FLOPS) for the proposed transformer neural operators, as implemented in practice, compared to FNO (Li et al., 2021). As done in Kaplan et al. (2020) and for ease of presentation, the additional evaluation cost arising from nonlinear activations, the softmax operator and normalizations are omitted. Furthermore, we omit the cost arising from skip connections. Indeed these do not affect the scaling. We further note that for our implementations, $d_{\textrm{FNO}}$, which is the latent dimension of the two-layer lifting and projection MLP in FNO, is fixed to be $128$. In this table we present the expression for the evaluation cost of AFNO without consideration of possible patching, which we do not employ for our comparisons: this in combination with the fact that we do not employ mode truncation is the reason for the $Nd_{\textrm{model}}^{2}$ term. In our context, the parameter $b$ in AFNO is set to $b=8$.